Joe can do a job twice as fast as Sam. Together they can do the job in 20 minutes. How long does it take Joe working alone to do the job?
Answer by AnlytcPhil(1806)

Joe can do a job twice as fast as Sam.
Let Sam's rate in jobs per hour be S.
Then Joe's rate in jobs per hour is 2S.

Together they can do the job in 20 minutes.
Their combined rate is 1 job per 20 minutes = 1 job per 1/3 hour or 3 jobs per hour.

So the sum of their rates is 3 jobs per hour

S + 2S = 3

    3S = 3

     S = 1

So Sam's rate is S = 1 job per hour, so it'll Sam 1 hour to do 1 job 

and Joe's rate is 2S = 2·1 = 2 jobs per hour, so it'll take Joe half
an hour (or 30 minutes) to do 1 job.

Edwin